# Towards Context-Aware Domain Generalization: Representing Environments with Permutation-Invariant Networks

### 15 Dec 2023 v1

@@ Abstract
In this work, we show that information about the context of an input $\mathbf{X}$ can improve the predictions of deep learning models when applied in new domains or production environments. We formalize the notion of context as a permutation-invariant representation of a set of data points that originate from the same environment/domain as the input itself. These representations are jointly learned with a standard supervised learning objective, providing incremental information about the unknown outcome. Furthermore, we offer a theoretical analysis of the conditions under which our approach can, in principle, yield benefits, and formulate two necessary criteria that can be easily verified in practice. Additionally, we contribute insights into the kind of distribution shifts for which our approach promises robustness. Our empirical evaluation demonstrates the effectiveness of our approach for both low-dimensional and high-dimensional data sets. Finally, we demonstrate that we can reliably detect scenarios where a model is tasked with unwarranted extrapolation in out-of-distribution (OOD) domains, identifying potential failure cases. Consequently, we showcase a method to select between the most predictive and the most robust model, circumventing the well-known trade-off between predictive performance and robustness.
@@ 1 Introduction
Figure 1: Conceptual sketch of our setup and approach. A) Data-generating process (DGP) that fulfills our criteria. We assume that the environment $E$ is a source node that is not caused by any system variable and that the relationship between $\mathbf{X}$ and $\mathbf{Y}$ varies with the environment. $\mathbf{S}^{(n)}$ is a set of $n$ IID inputs available in the new environment (i.e. context). B) The source component shift corresponding to the assumed DGP and example data coinciding with Simpson’s paradox. C) The workings of our approach in a test environment. A set-encoder generates a permutation-invariant representation $h(\mathbf{S}^{(n)})$ of the context. An inference network (e.g., a classifier) processes the representation along with the target input $\mathbf{X}$ and predicts the unknown outcome (e.g., label) of the target input. The set-representation can be combined with the input to reliably detect out-of-distribution (OOD) queries and prevent failure cases in domain generalization due to model misspecification.
Distribution shifts are the cause of many failure cases in machine learning [27, 32] and the root of various peculiar phenomena in classical statistics, such as Simpsons’ paradox [45, 57]. In this work, we employ permutation-invariant neural networks as set-encoders [15, 7] to improve the predictions of standard supervised models under distribution shift. Specifically, we consider training in the realm of Domain Generalization (DG), a setting where data from distinct environments [^1] We use the terms environment and domain interchangeably. is available for training and testing [39, 74].
@@
For illustration, consider a probabilistic model $p(\mathbf{Y}\mid\mathbf{X})$ that classifies diseases $\mathbf{Y}$ from magnetic resonance (MR) images $\mathbf{X}$. Since MR images are not fully standardized, the classifier should work slightly differently for images acquired by different hardware brands. It thus makes sense to inform the classifier about the current environment $E$ (here: hardware brand) and extend it into $p(\mathbf{Y}\mid\mathbf{X},E)$. This raises two questions:
  • (1)
    Under which circumstances will the classifier $p(\mathbf{Y}\mid\mathbf{X},E)$ be superior to $p(\mathbf{Y}\mid\mathbf{X})$?
  • (2)
    How should $E$ be represented to maximize the performance gain?
@@
The first question is important because there might exist a function $E=f(\mathbf{X})$ allowing the classifier $p(\mathbf{Y}\mid\mathbf{X})$ to deduce $E$ from the data $\mathbf{X}$. For example, $E$ might be inferred from the periphery of the given image, while $\mathbf{Y}$ depends on its central region. Then, no additional information is gained by passing $E$ explicitly, and both classifiers perform identically.
A straightforward answer to the second question is to distinguish environments by discrete labels. However, we argue that $E$ should be a continuous embedding. First, continuous embeddings can also be computed for new environments that have not been present in the training data. Second, when $p(\mathbf{Y}\mid\mathbf{X},E)$ receives a continuous $E$, it can potentially configure itself for unseen environments by interpolating between the training environments. And finally, discrete and continuous $E$ are equally informative for the known environments, ensuring no loss in information.
In the current work, we systematically investigate both questions, formalize three criteria when $p(\mathbf{Y}\mid\mathbf{X},E)$ is beneficial, and demonstrate how continuous embeddings $E$ can be learned from auxiliary data by means of set encoders (see Figure 1). Notably, two of these criteria are empirically testable using standard models and are shown to be necessary conditions for the success of the approach. First, we require that a single input alone is insufficient to deduce the originating environment (see Criterion 2.2). If this condition is not met, our model cannot possibly extract more information about the originating environment, compared to a standard model. Second, we require that, if the environment label is known by an oracle and given as an additional input to the standard model, its performance must improve (see Criterion 2.3). Otherwise, the set-based representation of the environment cannot possibly yield benefits. Notably, these two criteria are easy to verify and also met under the source component shift (see Section 2.6), which is a dataset shift that occurs in many scenarios [46].
@@
When test environments are highly dissimilar to the training environments, all DG methods enter an extrapolation regime with unknown prospects of success and the potential to generate silent failures. While our approach is not exempt from this “curse of extrapolation”, it comes with a natural way to reliably detect novel environments in set-representation space and delineate its competence region [41]. Moreover, we propose a method to select between models that are specialized in the in-distribution (ID) setting vs. models that are robust to out-of-distribution (OOD) scenarios on the fly. Thus, we can overcome the notorious trade-off between ID predictive performance and robustness to distribution shifts [65, 42, 37]. Accordingly, we can adaptively select the most robust model in the OOD setting and the most predictive model in the ID setting, an approach we demonstrate on the ColoredMNIST data set [4]. In summary, our contributions are:
  • We propose a novel approach to Domain Generalization (DG) that leverages context information from new environments in the form of learnable set-representations;
@@
  • We formalize the necessary and empirically verifiable conditions under which our approach can reap benefits from context information and improve on standard approaches;
@@
  • We perform an extensive empirical evaluation and show that we can reliably detect failure cases when the necessary criteria of our theory are not met, or when extrapolation is required.
@@ 2.1 Notation
We denote inputs $\mathbf{X}\in\mathcal{X}$ and outputs as $\mathbf{Y}\in\mathcal{Y}$, without any strict requirements on the input and output spaces $\mathcal{X}$ and $\mathcal{Y}$, respectively. We treat the (unknown) domain label $E$ as a random variable and denote with $\mathbf{S}^{(n)}$ an i.i.d. sample (i.e., a set of further inputs) from the given domain. The domain label $E$ is only known during training time and unknown during inference.
2.2 General Idea
The key idea of our approach is to build models that utilize not only a singleton input $\mathbf{X}$ to predict a target $\mathbf{Y}$, but also information about the environment of $\mathbf{X}$ that can improve the prediction $\hat{\mathbf{Y}}$. Providing environmental information in the form of a one-hot label is hardly feasible, as it presupposes the exact number of possible environments to be known during training, and that we always know from which environment an input originates at inference time. Consequently, such an approach is doomed to fail when the test data originates from a novel or unknown environment.
To overcome this problem, we employ permutation-invariant neural networks to adaptively represent environmental information given a set of test inputs. We will first detail our approach and then discuss criteria under which we can expect to reap benefits from the additional set-representation. Afterwards, we explain the theoretical data-generating process that matches these criteria, providing insight into the distribution shifts for which our approach may prove advantageous in practice. Finally, we discuss the process of identifying new environments that demand extrapolation, potentially leading to failure cases.
2.3 Permutation-Invariant Neural Networks
As mentioned above, a basic goal of our approach is to synthesize contextual information about a target input $\mathbf{X}$ by compressing a set of $n$ further inputs $\mathbf{S}^{(n)}:=\{\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{n}\}$ from the same environment into a permutation-invariant representation. The notion of permutation invariance is closely related to a core concept in probabilistic modeling and Bayesian inference – exchangeability [44]. Accordingly, an exchangeable sequence of random vectors is characterized by a joint distribution which is invariant to any permutation of the elements:
$p(\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{n})=p(\mathbf{X}_{\pi(1)},\mathbf{X}_{\pi(2)},\dots,\mathbf{X}_{\pi(n)}),$ (1)
where $\pi:\mathbb{N}\rightarrow\mathbb{N}$ denotes an arbitrary permutation of index elements $n\in\mathbb{N}$.
Exchangeable observations may come in various forms, for instance, patients entering a hospital, different measurements obtained with the same device, or sets of visual images, such as faces in a crowd. When it comes to learning exchangeable symmetries, permutation-invariant functions can serve as a key building block in neural architectures [7], as they automatically encode a favorable inductive bias towards permutation-invariance by design. A simple and intuitive way to create permutation-invariant functions involves the sum-decomposition
$h(\mathbf{S}^{(n)})=\rho\left(\sum_{i=1}^{n}\sigma(\mathbf{X}_{i})\right),$ (2)
where $\sigma$ and $\rho$ can be any functions, including deep neural networks [69, 7]. $h$ is permutation-invariant due to the summation operator which ensures that the argument of $\rho$ is agnostic to the order of elements in the set $\mathbf{S}^{(n)}$.
Despite having favorable theoretical properties, plain sum-decompositions can have limited representational capacity in practice [69, 58]. Thus, more expressive permutation-invariant functions can be learned by stacking equivariant transformations with sum-decompositions [69, 7] or by using a self-attention mechanism without positional encodings [33]. In order to re-assess the expressiveness of these methods, we show a comparison of the binary domain classification accuracy based on domain overlap and set size in Figure 10.
2.4 Context-Aware Model
Our model consists of two key components (also illustrated in Figure 1):
  • a permutation-invariant network $h_{\boldsymbol{\psi}}$ (“set encoder”) with parameters $\psi$ that maps a set input $\mathbf{S}^{(n)}$ to a summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, and
  • an inference network $f_{\boldsymbol{\phi}}$ with parameters $\boldsymbol{\phi}$ that maps both the input $\mathbf{X}$ and the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to a final prediction.
@@ 2.4 Context-Aware Model
The complete model is denoted as $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})=f_{\boldsymbol{\phi}}(\mathbf{X},h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)}))$ with parameters $\boldsymbol{\theta}=(\boldsymbol{\psi},\boldsymbol{\phi})$ for short. For a given supervised learning task, we aim to find the minimum to the following optimization problem
@@
$\boldsymbol{\theta}^{\star}=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathbb{E}_{p(\mathbf{X},\mathbf{Y},E)}\left[c(f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)}),\mathbf{Y})\right],$ (3)
where $c$ is a task-specific loss function (e.g., cross-entropy for classification or mean squared error for regression). The algorithmic details for optimizing Equation 3 are detailed in Algorithm 1. For practical reasons, we first apply a feature extractor $g$ and then pass $\{g(\mathbf{X})\mid\mathbf{X}\in\mathbf{S}^{(n)}\}$ as input to the set encoder and $g(\mathbf{X})$ as input to the inference network, building upon the features extracted by $g$. Note, that $g$ can be a pre-trained network, in which case we can treat it as a fixed transformation, or its parameters can be optimized jointly with $\boldsymbol{\theta}$.
Data: Samples from the joint distribution $p(\mathbf{X},\mathbf{Y},E)$
1 for $i=1,\dots,k$ do
2       Sample mini-batch =$\mathcal{B}=\{(\mathbf{x}_{1},\boldsymbol{y}_{1},\text{env}_{1}),\dots,(\mathbf{x}_{m},\boldsymbol{y}_{m},\text{env}_{m})\}$ from $p(\mathbf{X},\mathbf{Y},E)$
3       for $j=1,\dots,m$ do
4             Sample set $\mathbf{s}_{j}^{(n)}=\{\mathbf{x}_{1},\dots\mathbf{x}_{n}\}$ from $p(\mathbf{X}\mid E=\text{env}_{j})$
5             Replace $\text{env}_{j}$ with $\mathbf{s}_{j}^{(n)}$ in $\mathcal{B}$
6       end for
7      Update $\boldsymbol{\theta}$ using adaptive mini-batch gradient descent (or any variant):
8 end for
Output : Trained context-aware model $f_{\boldsymbol{\theta}}$
@@ 2.5 Criteria for Improvement
In the following, we establish criteria under which our method can exploit the distribution shifts between environments and yield improved predictions. In total, we propose three criteria that are necessary to achieve incremental improvement. In Theorem 2.1, we show how these criteria are related to each other. In the formulations below, $I(\mathbf{X};\mathbf{Y})$ denotes the mutual information between random vectors $\mathbf{X}$ and $\mathbf{Y}$ and $I(\mathbf{X};\mathbf{Y}\mid\mathbf{Z})$ denotes the conditional mutual information given a third random vector $\mathbf{Z}$. The symbol $\perp$ (resp. ⟂$\not\perp$̸) between two random vectors $\mathbf{X}$ and $\mathbf{Y}$ is used to express that the random vectors are independent (resp. dependent) or conditionally independent (resp. dependent) given a third random vector $\mathbf{Z}$.
@@
$\mathbf{S}^{(n)}\not\perp\mathbf{Y}\mid\mathbf{X}$ or equivalently $I(\mathbf{S}^{(n)};\,\mathbf{Y}\mid\mathbf{X})>0$.
@@
$E\not\perp\mathbf{S}^{(n)}\mid\mathbf{X}$ or equivalently $I(E;\,\mathbf{S}^{(n)}\mid\mathbf{X})>0$.
@@
The third criterion requires that the singleton input $\mathbf{X}$ gains information about $\mathbf{Y}$ if we also consider the origin environment $E$ of $\mathbf{X}$.
@@
$\mathbf{Y}\not\perp E\mid\mathbf{X}$ or equivalently $I(\mathbf{Y};\,E\mid\mathbf{X})>0$.
In Figure 1, this is evidently the case: If we knew the environment from which the data stems, we could improve our prediction of $\mathbf{Y}$. Furthermore, this criterion can serve as a sanity check in case we have an oracle that can identify the origin environment of the data with perfect accuracy.
In what follows, we show that Criterion 2.2 and Criterion 2.3 are necessary conditions for Criterion 2.1. We furthermore prove that if we can extract the environment label fully from $\mathbf{S}^{(n)}$, then Criterion 2.2 and Criterion 2.3 are sufficient conditions for Criterion 2.1. We even generalize this result for the case where the environment label is not inferable with 100% accuracy.
@@
  • (d)
The proof for this theorem and a discussion of its presumptions can be found in Section A.2. Unfortunately, we cannot conclude that $Y\perp\mathbf{S}^{(n)}\mid\mathbf{X}$ follows from Criterion 2.2 and Criterion 2.3 in general. An example where Criterion 2.2 and Criterion 2.3 hold, but Criterion 2.1 is violated, is provided in Section A.1.

### 21 Feb 2024 v2

# Towards Context-Aware Domain Generalization: Understanding the Benefits and Limits of Marginal Transfer Learning

@@ Abstract
In this work, we analyze the conditions under which information about the context of an input $\mathbf{X}$ can improve the predictions of deep learning models in new domains. Following work in marginal transfer learning in Domain Generalization (DG), we formalize the notion of context as a permutation-invariant representation of a set of data points that originate from the same domain as the input itself. We offer a theoretical analysis of the conditions under which this approach can, in principle, yield benefits, and formulate two necessary criteria that can be easily verified in practice. Additionally, we contribute insights into the kind of distribution shifts for which the marginal transfer learning approach promises robustness. Empirical analysis shows that our criteria are effective in discerning both favorable and unfavorable scenarios. Finally, we demonstrate that we can reliably detect scenarios where a model is tasked with unwarranted extrapolation in out-of-distribution (OOD) domains, identifying potential failure cases. Consequently, we showcase a method to select between the most predictive and the most robust model, circumventing the well-known trade-off between predictive performance and robustness.
@@ 1 Introduction
Figure 1: Conceptual sketch of our setup and approach. A) Data-generating process (DGP) that fulfills our criteria. We assume that the environment $E$ is a source node that is not caused by any system variable and that the relationship between $\mathbf{X}$ and $\mathbf{Y}$ varies with the environment. $\mathbf{S}^{(n)}$ is a set of $n$ i.i.d. inputs available in the new environment (i.e. context). B) The workings of the context-aware model (aka the marginal transfer learning approach) in a test environment. A set-encoder generates a permutation-invariant representation $h(\mathbf{S}^{(n)})$ of the context. An inference network (e.g., a classifier) processes the representation along with the target input $\mathbf{X}$ and predicts the unknown outcome (e.g., label) of the target input. The set-representation can be combined with the input to reliably detect out-of-distribution (OOD) queries and prevent failure cases in domain generalization due to model misspecification.
Distribution shifts are the cause of many failure cases in machine learning [Hendrycks and Dietterich, 2019, Koh et al., 2021] and the root of various peculiar phenomena in classical statistics, such as Simpson’s paradox [Peters et al., 2017, von Kügelgen et al., 2021]. The Domain Generalization (DG) task seeks a model that is robust to distribution shifts by utilizing data from distinct environments during training [Muandet et al., 2013, Zhou et al., 2022].
In the context of DG, marginal transfer learning enhances a model with context information to achieve better predictions Blanchard et al. [2021]. The “context” of a test instance is simply a set of samples that stems from the same environment as the instance itself and can be embedded, for instance, by permutation-invariant neural networks Edwards and Storkey [2017], Bloem-Reddy and Teh [2020]. In this work, we enhance the fundamental understanding of settings where marginal transfer learning in DG can reap benefits compared to baseline models.
@@
Consider a probabilistic model $p(\mathbf{Y}\mid\mathbf{X})$ that classifies diseases $\mathbf{Y}$ from magnetic resonance (MR) images $\mathbf{X}$. Since MR images are not fully standardized, the classifier should work slightly differently for images acquired by different hardware brands. It thus makes sense to inform the classifier about the current environment $E$ (here: hardware brand) and extend it into $p(\mathbf{Y}\mid\mathbf{X},E)$. This raises two questions: (i) Under which circumstances will the classifier $p(\mathbf{Y}\mid\mathbf{X},E)$ be superior to $p(\mathbf{Y}\mid\mathbf{X})$ and (ii) how should $E$ be represented to maximize the performance gain?
@@
The first question is important because there might exist a function $E=f(\mathbf{X})$ allowing the classifier $p(\mathbf{Y}\mid\mathbf{X})$ to deduce $E$ from the data $\mathbf{X}$ alone. For example, $E$ might be inferred from the periphery of the given image, while $\mathbf{Y}$ depends on its central region. Then, no additional information is gained by passing $E$ explicitly, and both classifiers perform identically. A straightforward answer to the second question is to distinguish environments by discrete labels, but we propose to learn continuous embeddings instead. This allows us to compute representations for novel environments that enable teal-time adaptation and interpolation between the training environments. Moreover, discrete and continuous $E$ are equally informative for the known environments, ensuring no loss in information.
Building on previous work in marginal transfer learning [Blanchard et al., 2021], we learn a continuous embedding of $E$ from auxiliary data using set encoders, as depicted in Figure 1. To systematically address the first question, we establish three criteria that delineate the circumstances in which $p(\mathbf{Y}\mid\mathbf{X},E)$ is beneficial, and subsequently prove their necessity. Notably, two of these criteria are empirically testable using standard models and are shown to be necessary conditions for the success of the approach.
@@
When test environments are highly dissimilar to the training environments, all DG methods enter an extrapolation regime with unknown prospects of success and an increased risk of silent failures. While marginal transfer learning is not exempt from this “curse of extrapolation”, we find that it comes with a natural way to reliably detect novel environments in set-representation space and delineate its competence region [Müller et al., 2023]. Accordingly, we propose a method to select between models that are specialized in the in-distribution (ID) setting vs. models that are robust to out-of-distribution (OOD) scenarios on the fly. Thus, we can overcome the notorious trade-off between ID predictive performance and robustness to distribution shifts [Yang et al., 2022b, Müller et al., 2022, Magliacane et al., 2018].
In summary, our contributions are:
@@
  • We formalize the necessary and empirically verifiable conditions under which the marginal transfer learning approach can reap benefits from context information and improve on standard approaches;
@@
  • We perform an extensive empirical evaluation and show that we can reliably detect failure cases when the necessary criteria of our theory are not met, or when extrapolation is required;
  • We show how the detection of novel environments allows for model selection, overcoming the trade-off between predictive performance and robustness.
@@ 2 Method
In the following, we will briefly introduce context-aware models (aka marginal transfer learning) in DG. We then discuss criteria under which we can expect to reap benefits from the additional set-representation. Afterwards, we explain the theoretical data-generating process that matches these criteria, providing insight into the distribution shifts for which context-aware approaches may prove advantageous in practice. Finally, we discuss the process of identifying new environments that demand extrapolation, potentially leading to failure cases.
@@ 2.1 Notation
We denote inputs $\mathbf{X}\in\mathcal{X}$ and outputs as $\mathbf{Y}\in\mathcal{Y}$, without any strict requirements on the input and output spaces $\mathcal{X}$ and $\mathcal{Y}$, respectively. We treat the (unknown) domain label $E$ as a random variable and denote with $\mathbf{S}^{(n)}$ a set of $n$ i.i.d. samples (i.e., a set of further inputs) from the given domain. The domain label $E$ is only known during training time and unknown during inference.
2.2 Context-Aware Models
@@
A context-aware model consists of two key components (also illustrated in Figure 1): (i) a permutation-invariant network $h_{\boldsymbol{\psi}}$ (“set encoder”) with parameters $\psi$ that maps a set input $\mathbf{S}^{(n)}$ to a summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, and (ii) an inference network $f_{\boldsymbol{\phi}}$ with parameters $\boldsymbol{\phi}$ that maps both the input $\mathbf{X}$ and the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to a final prediction. The complete model is denoted as $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})=f_{\boldsymbol{\phi}}(\mathbf{X},h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)}))$ with parameters $\boldsymbol{\theta}=(\boldsymbol{\psi},\boldsymbol{\phi})$ for short. For a given supervised learning task, we aim to find the minimum to the following optimization problem
@@
$\widehat{\boldsymbol{\theta}}=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathbb{E}_{p(\mathbf{X},\mathbf{Y},E)}\left[c(f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)}),\mathbf{Y})\right],$ (1)
where $c$ is a task-specific loss function (e.g., cross-entropy for classification or mean squared error for regression). Algorithm 1 details the optimization ofEquation 1.
@@ 2.3 Criteria for Improvement
In the following, we establish criteria under which context information allows to exploit the distribution shifts between environments and yield improved predictions. In total, we propose three criteria that are necessary to achieve incremental improvement. In Theorem 2.1, we show how these criteria are related to each other. In the formulations below, $I(\mathbf{X};\mathbf{Y})$ denotes the mutual information between random vectors $\mathbf{X}$ and $\mathbf{Y}$ and $I(\mathbf{X};\mathbf{Y}\mid\mathbf{Z})$ denotes the conditional mutual information given a third random vector $\mathbf{Z}$. The symbol $\perp$ (resp. ⟂$\not\perp$̸) between two random vectors $\mathbf{X}$ and $\mathbf{Y}$ is used to express that the random vectors are independent (resp. dependent) or conditionally independent (resp. dependent) given a third random vector $\mathbf{Z}$.
@@
$\mathbf{S}^{(n)}\not\perp\mathbf{Y}\mid\mathbf{X}$ or $I(\mathbf{S}^{(n)};\,\mathbf{Y}\mid\mathbf{X})>0$.
@@
$E\not\perp\mathbf{S}^{(n)}\mid\mathbf{X}$ or $I(E;\,\mathbf{S}^{(n)}\mid\mathbf{X})>0$.
@@
The third criterion requires that the singleton input $\mathbf{X}$ carries information about $\mathbf{Y}$ if we also consider the origin environment $E$ of $\mathbf{X}$.
@@
$\mathbf{Y}\not\perp E\mid\mathbf{X}$ or $I(\mathbf{Y};\,E\mid\mathbf{X})>0$.
This criterion can serve as a sanity check in case we have an oracle that can identify the origin environment of the data with perfect accuracy.
In what follows, we show that Criterion 2.2 and Criterion 2.3 are necessary conditions for Criterion 2.1. We furthermore prove that if we can extract the environment label fully from $\mathbf{S}^{(n)}$, then Criterion 2.2 and Criterion 2.3 are sufficient conditions for Criterion 2.1. We generalize this result for the case where the environment label is not inferable with 100% accuracy in Appendix C.
@@
The proof for this theorem can be found in subsection C.3. Unfortunately, we cannot conclude that $Y\perp\mathbf{S}^{(n)}\mid\mathbf{X}$ follows from Criterion 2.2 and Criterion 2.3 in general. A counterexample where Criterion 2.2 and Criterion 2.3 hold, but Criterion 2.1 is violated, is provided in subsection C.2. It is also worth noting that Criterion 2.3 might be achievable while Criterion 2.2 is unattainable and vice versa. For instance, when we can infer the originating environment from one sample (Criterion 2.2 is not attainable), the relation between $\mathbf{X}$ and $\mathbf{Y}$ might still vary with the environment (Criterion 2.3 is achievable).